Lauren is twice as old as her sister, Sue. The sum of their ages is 24. How old are the sisters ?
Answer by lenny460(1073)

Lauren is twice as old as her sister, Sue. The sum of their ages is 24. How old are the sisters ?

Let the age of Sue be n
Let the age of Lauren be 2n
Therefore:
n + 2n = 24
n + 2n = 3n
3n = 24
Divide each side by 3
3n/3 = 24/3
n = 8
2n = 2 * 8 = 16


Answer:
Sue is 8 years old and Lauren is 16 years old.









Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com